# Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf

### v1

@@ C.1 Examples of retrieved context
Fig. 9 displays examples of context for RANa ObjectNav ([Uncaptioned image]) using similarity in CLIP feature space for the six different goal categories, chair, bed, plant,sofa, tv monitor and toilet. As discussed in the paper, results accurately represent instances of the target object present in the scene, with the exception of the tv monitor class where we observe some errors, especially in scenes where monitors are absent.
@@ C.2 Navigation examples
In Fig. 10 we show examples of our agent RANa-b navigating to two challenging ImageNav goals. The figures visualize the observation (left), goal (highlighted in green), $N=8$ context elements, with the one selected at the given time step highlighted in red, and the agent path drawn on the scene floorplan on the right. Each row of pictures focuses on one time step along the episode and are ordered in time from top to bottom. In this figure and the following ones, time is not sampled uniformly: we selected time steps, typically towards the end of the episode, where the agent makes interesting context selections to solve the episode.
The episode in Fig. 10 (left) is challenging because the environment is rather complex, the goal image is ambiguous and it is “hidden” in a terrace at the extreme perifery of the navigable space. The baseline DEBiT-b agent, without context, does its best to explore the apartment and gets close to the terrace door where the goal is located, but does not manage to navigate to it. RANa-b selects from the context an image (highlighted in red) that contains the railing of the terrace that helps our agent solve this complex navigation episode. The navigation episode on the right is complex for similar reasons: large complex environment and goal image with very little information. RANa is capable of exploiting the few relevant context items, by first selecting the context image containing the door to the room where the goal is located, and then the cabinet, whose corner is visible on the goal image.
Figure 11: Refer to caption  RANa-b successfully navigating to two goals in zero-shot Instance-ImageNav episodes - The agent observation is on the left, next to it is the original goal image from the Instance-ImageNav task (in cyan) and the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ highlighted in green. The remaining data is organized and visualized as in Fig. 10.
Fig. 11 shows RANa-b ImageNav agent successfully navigating in two zero-shot Instance-ImageNav episodes. Results are visualized in the same way as in Fig. 10, with the exception of the goal image. In this case we have in fact two goal images, one from the original task (highlighted in cyan) which is taken with a sensor with different pose and intrinsics of the agent sensor, and the goal image $\mathbf{g}^{\mathcal{R}}$, highlighted in green, which is retrieved from the scene database $\mathcal{R}$. In both cases, RANa selects context items that represent elements visible in our agent observation (the toilet on the left and the door on the right) that guide the agent inside the room where the goal is located. Interestingly, in both cases the baseline agent passed along the same path, but did not enter the room, arguably because it did not have access to the context.
@@
In Fig. 12 we present an even more challenging scenario, where RANa navigates in two zero-shot Instance-ImageNav episodes, where the goal image is captured with a smartphone in a real environment. As mentioned in the main text, we have access to this environment and a 3D model of the scene is available. The agent navigates to the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ in the simulated reconstruction of the scene. The example on the left is particularly interesting because the retrieved goal image, in green, is not the correct instance of the original goal. However, guided by several context items that contain instances of the original goal image, RANa succeeds in reaching the goal. In the navigation episode on the right, our agent picks relevant context elements containing the chairs of the meeting room where the goal is positioned, and enters the room solving the navigation episode.
@@
Finally, we observe cases in which access to a context deteriorates the performance of the baseline agent. We can describe the behavior of our RANa agent in these cases as being “distracted” by context items visible by the agent at a given time step. Fig. 13 shows a typical instance of this phenomenon: the baseline DEBiT agent (trajectory shown at the bottom) quickly gets to the goal, while RANa reaches the goal area, but starts turning on itself, cycling multiple times through relevant, but “distracting”, context images.

### v2

@@ E.1 Examples of retrieved context for ObjectNav
Fig.˜8 displays examples of context for RANa ObjectNav ([Uncaptioned image]) using similarity in CLIP feature space for the six different goal categories, chair, bed, plant, sofa, tv monitor and toilet. Results accurately represent instances of the target object present in the scene, with the exception of the tv monitor class where we observe some errors, especially in scenes where monitors are absent.
@@ E.2 Navigation examples
In Fig.˜9 we show examples of our agent RANa-b static navigating to two challenging ImageNav goals. The figures visualize the observation (left), goal (highlighted in green), $N=8$ context elements, with the one selected at the given time step highlighted in red, and the agent path drawn on the scene floorplan on the right. Each row of pictures focuses on one time step along the episode and are ordered in time from top to bottom. In this figure and the following ones, time is not sampled uniformly: we selected time steps, typically towards the end of the episode, where the agent makes interesting context selections to solve the episode.
The episode in Fig.˜9 (left) is challenging because the environment is rather complex, the goal image is ambiguous and it is “hidden” in a terrace at the extreme perifery of the navigable space. The baseline DEBiT-b agent, without context, does its best to explore the apartment and gets close to the terrace door where the goal is located, but does not manage to navigate to it. RANa-b selects from the context an image (highlighted in red) that contains the railing of the terrace that helps our agent solve this complex navigation episode. The navigation episode on the right is complex for similar reasons: large complex environment and goal image with very little information. RANa is capable of exploiting the few relevant context items, by first selecting the context image containing the door to the room where the goal is located, and then the cabinet, whose corner is visible on the goal image.
Figure 10: Refer to caption  RANa-b successfully navigating to two goals in zero-shot Instance-ImageNav episodes - The agent observation is on the left, next to it is the original goal image from the Instance-ImageNav task (in cyan) and the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ highlighted in green. The remaining data is organized and visualized as in Fig.˜9.
Fig.˜10 shows RANa-b ImageNav agent successfully navigating in two zero-shot Instance-ImageNav episodes. Results are visualized in the same way as in Fig.˜9, with the exception of the goal image. In this case we have in fact two goal images, one from the original task (highlighted in cyan) which is taken with a sensor with different pose and intrinsics of the agent sensor, and the goal image $\mathbf{g}^{\mathcal{R}}$, highlighted in green, which is retrieved from the scene database $\mathcal{R}$. In both cases, RANa selects context items that represent elements visible in our agent observation (the toilet on the left and the door on the right) that guide the agent inside the room where the goal is located. Interestingly, in both cases the baseline agent passed along the same path, but did not enter the room, arguably because it did not have access to the context.
@@
In Fig.˜11 we present an even more challenging scenario, where RANa navigates in two zero-shot Instance-ImageNav episodes, where the goal image is captured with a smartphone in a real environment. As mentioned in the main text, we assume having access to this environment and that a 3D model of the scene is available. The agent navigates to the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ in the simulated reconstruction of the scene. The example on the left is particularly interesting because the retrieved goal image, in green, is not the correct instance of the original goal. However, guided by several context items that contain instances of the original goal image, RANa succeeds in reaching the goal. In the navigation episode on the right, our agent picks relevant context elements containing the chairs of the meeting room where the goal is positioned, and enters the room solving the navigation episode.
@@
Finally, we observe cases in which access to a context deteriorates the performance of the baseline agent. We can describe the behavior of our RANa agent in these cases as being “distracted” by context items visible by the agent at a given time step. Fig.˜12 shows a typical instance of this phenomenon: the baseline DEBiT agent (trajectory shown at the bottom) quickly gets to the goal, while RANa reaches the goal area, but starts turning on itself, cycling multiple times through relevant, but “distracting”, context images.